# Derivative-enhanced Deep Operator Network

### v1

@@ A.1 Proof of the Equivalence of Equation 6 and Equation 7
A.2 Proof of the Equivalence of Equation 8 and Equation 10
@@ Appendix B Experimental Details
B.1 Data generation
@@
B.2 Training details
@@
For the vanilla DeepONet, we parameterize the branch net using a CNN and the trunk net using a ResNet. For the reduced basis DeepONet, both the branch net and trunk net are parameterized using ResNets. We train each model for 1000 epochs. See Table 5, Table 6 and Table 7 for details about other hyperparameters we have used. When the loss functions comprises more than two type of losses, we enhance the model’s performance and training stability by incorporating the self-adaptive learning rate annealing algorithm [19], with an update frequency of $100$ and a moving average parameter of $0.9$. This algorithm automatically adjusts the loss weights $\{\lambda_{i}\}_{i=1}^{3}$ in Equation 5 at a minimal computational cost. Additionally, we standardize the inputs and labels before training.
Table 5: Training details for the vanilla DeepONet
@@
Darcy flow Hyperelasticity Navier–Stokes
Branch net CNN (7 layers, 128 out dim, ReLU) CNN (7 layers, 128 out dim, ReLU) CNN (7 layers, 128 out dim, ReLU)
Trunk net ResNet (7 layers, 128 out dim, ReLU) ResNet (7 layers, 256 out dim, ReLU) ResNet (7 layers, 256 out dim, ReLU)
Initialization Kaiming Uniform Kaiming Uniform Kaiming Uniform
(lr, weight decay) $(0.001,10^{-6})$ $(0.001,10^{-6})$ $(0.001,10^{-5})$
(gamma, step size) $(0.5,10)$ $(0.6,10)$ $(0.7,10)$
Table 6: Training details for the DE-DeepONet (KLE basis + dm labels + dx labels)
@@
Darcy Hyperelasticity Navier–Stokes
Branch net ResNet (4 layers, 128 out dim, ELU) ResNet (4 layers, 128 out dim, ELU) ResNet (7 layers, 128 out dim, ELU)
Trunk net ResNet (4 layers, 128 out dim, ELU) ResNet (4 layers, 256 out dim, ELU) ResNet (7 layers, 256 out dim, ELU)
Initialization Kaiming Normal Kaiming Normal Kaiming Normal
(lr, weight decay) $(0.005,10^{-13})$ $(0.001,10^{-12})$ $(0.001,0)$
(gamma, step size) $(0.9,10)$ $(0.99,50)$ $(0.9,20)$
$N_{x}^{\text{batch}}(=\alpha N_{x})$ $633\approx(0.15\times 65^{2})$ $422\approx(0.1\times 65^{2})$ $422\approx(0.1\times 65^{2})$
Table 7: Training details for the DE-DeepONet (ASM basis + dm labels + dx labels)
@@
Darcy Hyperelasticity Navier–Stokes
Branch net ResNet (4 layers, 128 out dim, ELU) ResNet (4 layers, 128 out dim, ELU) ResNet (7 layers, 128 out dim, ELU)
Trunk net ResNet (4 layers, 128 out dim, ELU) ResNet (4 layers, 256 out dim, ELU) ResNet (7 layers, 256 out dim, ELU)
Initialization Kaiming Normal Kaiming Normal Kaiming Normal
(lr, weight decay) $(0.005,10^{-13})$ $(0.001,10^{-12})$ $(0.001,10^{-12})$
(gamma, step size) $(0.9,10)$ $(0.99,50)$ $(0.9,10)$
$N_{x}^{\text{batch}}(=\alpha N_{x})$ $633\approx(0.15\times 65^{2})$ $422\approx(0.1\times 65^{2})$ $422\approx(0.1\times 65^{2})$
B.3 Test errors
We present the comparisons of relative error on test dataset (using the metrics in Section 4.3) of different methods considered in this work in Tables 8, LABEL:, 9, LABEL: and 10 for the Darcy flow equation, Tables 11, LABEL:, 12, LABEL: and 13 for the hyperelasticity equation, Tables 14, LABEL:, 15, LABEL: and 16 for the Navier–Stokes equations. For visualization, we provide the mean relative error plots in Figure 8 for the Darcy flow equation, Figure 9 for the hyperelasticity equation, and Figures 10, LABEL: and 4 for the Navier–Stokes equations.
Table 8: Darcy $5$ trails test relative $L^{2}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 48.19 (5.53) 23.86 (0.77) 15.31 (0.75) 6.28 (0.33)
KLE 67.13 (4.94) 27.51 (0.95) 11.06 (0.15) 7.66 (0.09)
KLE + dm 24.40 (1.50) 10.46 (0.37) 4.89 (0.04) 4.44 (0.01)
KLE + dm + dx 22.34 (1.08) 9.39 (0.27) 4.87 (0.04) 4.47 (0.01)
ASM 63.78 (4.01) 26.26 (0.76) 10.76 (0.33) 6.87 (0.09)
ASM + dm 24.54 (1.26) 8.95 (0.18) 4.54 (0.03) 4.06 (0.02)
ASM + dm + dx 22.12 (1.03) 8.08 (0.24) 4.57 (0.02) 4.12 (0.01)
Table 9: Darcy $5$ trails test relative $H^{1}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 83.11 (3.89) 50.02 (1.93) 35.34 (1.12) 20.69 (0.36)
KLE 85.88 (7.92) 44.78 (1.22) 24.75 (0.23) 21.60 (0.12)
KLE + dm 48.27 (1.65) 27.16 (0.74) 17.44 (0.15) 15.98 (0.07)
KLE + dm + dx 39.11 (1.67) 22.57 (0.31) 16.38 (0.08) 15.41 (0.03)
ASM 82.75 (5.79) 42.62 (1.10) 24.22 (0.25) 20.43 (0.19)
ASM + dm 41.07 (1.30) 24.59 (0.33) 17.03 (0.09) 15.67 (0.08)
ASM + dm + dx 35.62 (0.45) 21.27 (0.25) 16.06 (0.07) 15.06 (0.04)
Table 10: Darcy $5$ trails test relative $Fro$ error (for dm): mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 86.27 (3.25) 61.71 (2.02) 45.89 (2.89) 25.89 (2.32)
KLE 119.02 (5.27) 65.16 (2.02) 29.90 (0.38) 23.18 (0.27)
KLE + dm 57.11 (1.82) 28.97 (1.09) 11.06 (0.20) 8.89 (0.03)
KLE + dm + dx 55.32 (1.81) 26.42 (0.85) 11.21 (0.19) 9.20 (0.04)
ASM 120.13 (6.00) 62.25 (2.26) 28.83 (0.59) 20.44 (0.14)
ASM + dm 61.02 (2.91) 24.45 (0.62) 10.68 (0.06) 8.82 (0.05)
ASM + dm + dx 56.74 (1.05) 22.59 (0.90) 10.86 (0.13) 9.20 (0.05)
@@ B.3 Test errors
Figure 8: Mean relative errors over 5 trials with a varying number of training samples for the Darcy equation using different methods. Relative errors in the $L^{2}(\Omega)$ norm (left) and $H^{1}({\Omega})$ norm (middle) for the prediction of $u$. Right: Relative error in the Frobenius (Fro) norm for the prediction of $du(m;\omega_{1})$.
Table 11: Hyperelasticity $5$ trails test relative $L^{2}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 45.89 (1.15) 33.28 (2.10) 17.64 (0.90) 7.82 (0.49)
KLE 80.54 (10.88) 29.29 (0.84) 12.81 (0.39) 9.14 (0.49)
KLE + dm 17.18 (1.23) 9.14 (0.76) 5.00 (0.13) 4.73 (0.34)
KLE + dm + dx 17.11 (1.50) 8.61 (0.60) 5.07 (0.21) 4.57 (0.26)
ASM 80.36 (6.16) 29.74 (3.25) 13.11 (1.15) 7.06 (0.23)
ASM + dm 21.81 (1.53) 8.71 (0.82) 4.61 (0.31) 3.47 (0.23)
ASM + dm + dx 20.24 (1.38) 8.58 (0.81) 4.90 (0.81) 3.61 (0.21)
Table 12: Hyperelasticity $5$ trails test relative $H^{1}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 57.80 (1.50) 41.87 (1.85) 26.23 (0.85) 16.77 (0.26)
KLE 80.91 (7.29) 38.88 (0.71) 21.73 (0.26) 17.46 (0.22)
KLE + dm 27.63 (0.91) 18.68 (0.17) 14.80 (0.13) 12.72 (0.27)
KLE + dm + dx 24.80 (1.01) 16.66 (0.49) 13.38 (0.31) 12.06 (0.26)
ASM 81.64 (5.59) 39.09 (1.88) 21.81 (0.50) 16.51 (0.15)
ASM + dm 31.15 (0.64) 18.76 (0.55) 14.56 (0.36) 13.02 (0.13)
ASM + dm + dx 27.38 (0.63) 16.64 (0.52) 13.26 (0.21) 11.92 (0.25)
Table 13: Hyperelasticity $5$ trails test relative $Fro$ error (for dm): mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 102.82 (0.86) 85.17 (5.59) 57.52 (3.76) 32.34 (2.33)
KLE 142.92 (11.76) 81.00 (3.57) 39.59 (0.74) 28.12 (0.83)
KLE + dm 50.93 (2.13) 25.82 (0.54) 15.39 (0.31) 13.08 (0.16)
KLE + dm + dx 43.09 (2.76) 22.16 (0.26) 12.93 (0.24) 10.91 (0.09)
ASM 138.26 (12.76) 81.01 (3.22) 37.78 (1.16) 22.47 (0.24)
ASM + dm 59.06 (1.67) 25.36 (1.41) 12.38 (0.29) 9.58 (0.05)
ASM + dm + dx 50.22 (2.48) 21.37 (0.84) 10.59 (0.12) 8.37 (0.20)
@@
Figure 9: Mean relative errors over 5 trials with a varying number of training samples for the hyperelasticity equation using different methods. Relative errors in the $L^{2}(\Omega)$ norm (left) and $H^{1}({\Omega})$ norm (middle) for the prediction of $u=(u_{1},u_{2})$. Right: Relative error in the Frobenius (Fro) norm for the prediction of $du(m;\omega_{1})=(du_{1}(m;\omega_{1}),du_{2}(m;\omega_{2}))$.
Table 14: Navier–Stokes $5$ trails test relative $L^{2}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 41.28 (1.17) 26.64 (1.34) 16.18 (0.86) 7.90 (0.86)
KLE 64.44 (3.01) 28.19 (1.27) 16.39 (0.09) 14.87 (0.12)
KLE + dm 24.04 (1.13) 13.84 (0.25) 11.06 (0.09) 10.44 (0.04)
KLE + dm + dx 21.89 (0.80) 13.52 (0.29) 11.06 (0.05) 10.45 (0.03)
ASM 41.61 (1.12) 18.31 (1.03) 11.16 (0.54) 9.27 (0.21)
ASM + dm 21.33 (0.62) 10.16 (0.24) 6.81 (0.05) 6.25 (0.01)
ASM + dm + dx 21.88 (0.66) 10.69 (0.25) 7.09 (0.04) 6.27 (0.01)
Table 15: Navier–Stokes $5$ trails test relative $H^{1}(\Omega)$ error: mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 66.18 (1.42) 47.20 (2.02) 23.10 (0.80) 13.17 (0.58)
KLE 61.97 (1.44) 34.53 (1.12) 21.86 (0.19) 21.45 (0.10)
KLE + dm 51.81 (2.07) 31.57 (1.11) 18.29 (0.27) 15.40 (0.02)
KLE + dm + dx 35.54 (0.91) 27.28 (0.23) 18.53 (0.12) 15.42 (0.03)
ASM 51.19 (2.65) 28.48 (2.90) 16.48 (1.70) 14.10 (0.29)
ASM + dm 48.08 (1.27) 26.71 (0.54) 13.28 (0.37) 10.57 (0.01)
ASM + dm + dx 47.78 (2.11) 27.28 (0.41) 17.87 (0.48) 10.91 (0.17)
Table 16: Navier–Stokes $5$ trails test relative $Fro$ error (for dm): mean(standard deviation%)
Method Number of training samples
16 64 256 1024
Vanilla 576.93 (42.09) 375.72 (51.98) 257.51 (25.55) 212.01 (42.70)
KLE 607.06 (18.07) 293.72 (15.57) 144.20 (4.31) 165.90 (4.86)
KLE + dm 199.44 (7.48) 113.26 (7.47) 68.60 (2.01) 56.65 (0.62)
KLE + dm + dx 197.82 (10.62) 113.70 (5.85) 69.96 (2.14) 58.49 (0.39)
ASM 195.21 (9.51) 78.99 (2.44) 50.55 (0.69) 49.05 (0.60)
ASM + dm 104.05 (1.11) 52.88 (1.51) 30.67 (0.16) 26.36 (0.08)
ASM + dm + dx 102.07 (1.80) 54.23 (1.54) 31.47 (0.18) 26.03 (0.07)
@@
Figure 10: Mean relative errors over 5 trials with a varying number of training samples for the Navier–Stokes equations using different methods. Relative errors in the $L^{2}(\Omega)$ norm (left) and $H^{1}({\Omega})$ norm (middle) for the prediction of the velocity-x. Right: Relative error in the Frobenius (Fro) norm for the prediction of the derivative of the velocity-x with respect to the parameter in the direction of the 128 samples drawn from the same distribution as parameter.
B.4 Reduced basis
We plot the eigenvalues of KLE basis and ASM basis in Figure 11. We present visualizations of the first three KLE and ASM bases in Figure 12.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 11: First 100 eigenvalues of KLE and ASM basis. Top row: Darcy; Middle row: Hyperelasticity; Bottom row: Navier–Stokes
@@ B.4 Reduced basis
Refer to caption Refer to caption Refer to caption
Figure 12: First row: Darcy first $3$ KLE basis (similar patterns observed for Hyperelasticity and Navier–Stokes); Second row: Darcy first $3$ ASM basis; Third row: Hyperelasticity first $3$ ASM basis; Fourth row: Navier–Stokes first $3$ ASM basis.

### v2

@@ A.2 Proof of the equivalence of Equation 6 and Equation 7
A.3 Proof of the equivalence of Equation 8 and Equation 10
@@ Appendix B Experimental details
B.1 Governing equations
Hyperelasticity equation.
$d\widetilde{W}(u;v)=0,$
holds for any test function $v$ in the state space. See Figure 3 for the visualization for one parameter-solution pair of the hyperelasticity equation.
Figure 3: Visualization of one parameter-solution pair of hyperelasticity equation. The color of the output indicates the magnitude of the displacement $u$ (which maps from domain $\Omega$ to $\mathbb{R}^{2}$) instead of its componentwise function $u_{1}$ or $u_{2}$. The skewed square shows locations of any domain point after deformation $X\to x$. See Figures 13 and 14 for $u_{1}$ and $u_{2}$.
Navier–Stokes equations.
where $\Gamma_{\text{top}}$ and $\Gamma$ denote the left and whole boundary of the cavity domain $\Omega$, respectively. Here, we assume that the randomness arises from the viscosity term $e^{m}$. Our goal is learn the operator that maps the parameter $m$ to the velocity $u$. For demonstration, we choose $\bar{m}=6.908$ ($e^{\bar{m}}\approx 10^{3}$, thus the viscosity term dominates), $\delta=0.6$, and $\gamma=0.06$. See Figure 4 for the visualization for one parameter-solution pair of the Navier–Stokes equations.
Figure 4: Visualization of one parameter-solution pair of Navier–Stokes equations.
B.2 Data generation
@@
B.3 Computation of derivative labels and outputs
B.3.1 Derivative labels $p:=du(m;\psi)$ as ground truth
Since the PDE residual $\mathcal{R}(m,u(m))\equiv 0$ holds for any $m\in V_{h}^{\text{in}}$, we have
$\mathcal{R}(m+\varepsilon\psi,u(m+\varepsilon\psi))=0,\quad\forall\varepsilon>0,\psi\in V_{h}^{\text{in}}.$
By the Taylor expansion and $\mathcal{R}(m,u(m))=0$, we obtain
where $\delta u=u(m+\varepsilon\psi)-u(m)$. Dividing both sides of Eq. (32) by $\varepsilon$ and letting $\varepsilon$ approach 00 yields
$(\partial_{m}\mathcal{R})\psi+(\partial_{u}\mathcal{R})p=0,$ (33)
where $p:=du(m;\psi)$. We solve Eq. (33) for $p$ via its weak form
$\langle(\partial_{m}\mathcal{R})\psi,v\rangle+\langle(\partial_{u}\mathcal{R})p,v\rangle=0,$
where $v$ is a test function in $V_{h}^{\text{out}}$.
Example.
Consider the (nonlinear) diffusion-reaction equation
$\begin{cases}&-\nabla\cdot(e^{m}\nabla u)+u^{3}=1,\quad x\in\Omega,\\ &u(x)=0,\quad x\in\partial\Omega.\end{cases}$
Then
  • $\mathcal{R}=-\nabla\cdot(e^{m}\nabla u)+u^{3}-1$
  • $(\partial_{m}\mathcal{R})\psi=-\nabla\cdot(e^{m}\psi\nabla u)$
  • $(\partial_{u}\mathcal{R})p=-\nabla\cdot(e^{m}\nabla p)+3u^{2}p$
Thus, $p$ satisfies the linear PDE
$\langle e^{m}\psi\nabla u,\nabla v\rangle+\langle e^{m}\nabla p,\nabla v\rangle+\langle 3u^{2}p,v\rangle=0.$
Using FEniCS, we can easily derive Gâteaux derivative via automatic symbolic differentiation instead of by hand. In this case, the Python code for representing the weak form of the residual $\langle\mathcal{R},v\rangle$ and Gâteaux derivatives $\langle(\partial_{m}\mathcal{R})\psi,v\rangle$ and $\langle(\partial_{u}\mathcal{R})p,v\rangle$ can be written as
import dolfin as dl
  • R=(dl.inner(dl.exp(m)*dl.grad(u), dl.grad(v))*dl.dx
    +(u**3-dl.Constant(1.0))*v*dl.dx)
  • dl.derivative(R,m,psi)
  • dl.derivative(R,u,p)
B.3.2 Derivative outputs of neural networks
DE-DeepONet.
For notation simplicity, we illustrate the case where the input reduced basis is ASM basis and the output function is real-valued. The output of the model is given by
$\hat{u}(\bm{m};\theta)(x)=\langle b((\Psi^{\text{in}})^{T}C^{-1}\bm{m};\theta_{b}),t(x;\theta_{t})\rangle+\theta_{\text{bias}},$
where $\Psi^{\text{in}}=(\bm{\psi}_{1}^{\text{in}}|\cdots|\bm{\psi}_{r_{\text{in}}}^{\text{in}})\in\mathbb{R}^{N_{h}^{\text{in}}\times r_{\text{in}}}$ are the nodal values of input (ASM) reduced basis functions, $C^{-1}$ is the inverse of the covariance matrix of the Gaussian random field $\nu$ where parameter $m$ is sampled from (Recall that the ASM basis are orthonormal in the inner product with weight matrix ${C}^{-1}$.) Thus, by the chain rule of derivative, for any test direction $\bm{\psi}_{\text{test}}$, one has
$\nabla_{\bm{m}}\hat{u}(\bm{m};\theta)(x)\bm{\psi}_{\text{test}}=\nabla_{\widetilde{m}}\langle b(\widetilde{m}),t(x;\theta_{t})\rangle(\Psi^{\text{in}})^{T}C^{-1}\bm{\psi}_{\text{test}},$
where $\widetilde{m}=(\Psi^{\text{in}})^{T}C^{-1}\bm{{m}}$. The Jacobian matrix $\nabla_{\widetilde{m}}\langle b(\widetilde{m}),t(x;\theta_{t})\rangle$ can be efficiently computed using, e.g., torch.func.jacrev, and further parallelized with torch.vmap. If $\bm{\psi}_{\text{test}}$ is in $\Psi^{\text{in}}$ during model training, we can see that $(\Psi^{\text{in}})^{T}C^{-1}\bm{\psi}_{\text{test}}$ becomes a unit vector which frees us the need to compute it, otherwise in the model inference stage when $\bm{\psi}_{\text{test}}$ is (the nodal values of) a random function sampled from $\nu$, we compute $T=(\Psi^{\text{in}})^{T}C^{-1}$ and then $T\bm{\psi}_{\text{test}}$.
DINO.
The output of the model is given by
$\hat{u}(\bm{m};\theta)(x)=\Phi^{\text{out}}(x)\Psi^{\text{out}}f_{\theta}((\Psi^{\text{in}})^{T}C^{-1}\bm{{m}}),$
where $\Phi^{\text{out}}(x)=(\phi_{1}^{\text{out}}(x),\ldots,\phi_{N_{h}^{\text{out}}}^{\text{out}}(x))\in\mathbb{R}^{1\times N_{h}^{\text{out}}}$ denotes the output finite element basis functions evaluated on point $x\in\Omega$, $\Psi^{\text{out}}=(\bm{\psi}_{1}^{\text{out}}|\cdots|\bm{\psi}_{r_{\text{out}}}^{\text{out}})\in\mathbb{R}^{N_{h}^{\text{out}}\times r_{\text{out}}}$ are the nodal values of output (POD) reduced basis functions, similarly for $\Psi^{\text{in}}$ denoting the input (ASM) reduced basis, and $C^{-1}$ is the inverse of the covariance matrix of the Gaussian random field where parameter $m$ is sampled from. Thus, by the chain rule of derivative, for any test direction $\bm{\psi}_{\text{test}}$, one has
$\nabla_{\bm{m}}\hat{u}(\bm{m};\theta)(x)\bm{\psi}_{\text{test}}=\Phi^{\text{out}}(x)\Psi^{\text{out}}\nabla_{\widetilde{m}}f_{\theta}(\widetilde{m})(\Psi^{\text{in}})^{T}C^{-1}\bm{\psi}_{\text{test}},$
where $\widetilde{m}=(\Psi^{\text{in}})^{T}C^{-1}\bm{{m}}$. For fast evaluations given different $x$, $\bm{m}$, and $\bm{\psi}_{\text{test}}$, we first compute $T_{\text{left}}=\Phi^{\text{out}}(x)\Psi^{\text{out}}$ for all points $x$ that need to be evaluated and $T_{\text{right}}=(\Psi^{\text{in}})^{T}C^{-1}$. Next, we compute $J=\nabla_{\widetilde{m}}f_{\theta}(\widetilde{m})$ using, e.g., torch.func.jacrev, and finally compute $T_{\text{left}}JT_{\text{right}}\bm{\psi}_{\text{test}}$.
DeepONet & FNO.
Both models receive the full high dimensional parameter $\bm{m}$, so we compute the directional derivative $\nabla_{\bm{m}}\hat{u}(\bm{m};\theta)\bm{\psi}_{\text{test}}$ using the Jacobian-vector product torch.jvp instead of the full Jacobian in DE-DeepONet and DINO. For the coordinates $x$ as additional inputs, we pad zeros to the tensor $\bm{\psi}_{\text{test}}$ to match the dimension of the input tensor $(\bm{m},\{x^{(j)}\}_{j=1}^{N_{x}})$.
B.4 Training details
@@
For the DeepONet, we parameterize the branch net using a CNN and the trunk net using a ResNet. For the FNO, we use the package neuraloperator. For the DINO, we use $16$ ASM basis functions for the input dimension reduction and $64$ POD basis functions for the output dimension reduction, and parameterize the neural network using a ResNet. For the DE-DeepONet, both the branch net and trunk net are parameterized using ResNets. We train each model for 32768 iterations (with the same batch size 8) using an AdamW optimizer [34] and a StepLR learning rate scheduler (We disable learning rate scheduler for DE-DeepONet). See Tables 7, 8, 9 and 10 for details. When the loss function comprises two terms, we apply the self-adaptive learning rate annealing algorithm from [23], with an update frequency of $100$ and a moving average parameter of $0.9$, to automatically adjust the loss weights $\{\lambda_{i}\}_{i=1}^{2}$ in Equation 5. Additionally, we standardize the inputs and labels before training.
Table 7: Training details for DeepONet
@@
Hyperelasticity Navier–Stokes
branch net CNN
  6 hidden layers
  256 output dim
  ReLU
CNN
  6 hidden layers
  256 output dim
  ReLU
trunk net ResNet
  3 hidden layers
  512 output dim
  512 hidden dim
  ReLU
ResNet
  3 hidden layers
  512 output dim
  512 hidden dim
  ReLU
initialization Kaiming Uniform Kaiming Uniform
AdamW (lr, weight decay) $(10^{-3},10^{-6})$ $(10^{-3},10^{-5})$
StepLR (gamma, step size) $(0.6,10)$ $(0.7,10)$
number of Fourier features $64$ $64$
Fourier feature scale $\sigma$ $0.5$ $1.0$
Table 8: Training details for FNO & DE-FNO
@@
Hyperelasticity Navier–Stokes
number of modes (32,32) $(32,32)$
in channels $3$ $3$
out channels $2$ $2$
hidden channels $64$ $64$
number of layers $4$ $4$
lifting channel ratio $2$ $2$
projection channel ratio $2$ $2$
activation function GELU GELU
AdamW (lr, weight decay) $(5\times 10^{-3},10^{-4})$ $(5\times 10^{-3},10^{-4})$
StepLR (gamma, step size) $(0.9,25)$ $(0.9,50)$
Table 9: Training details for DINO
@@
Hyperelasticity Navier–Stokes
neural network ResNet
  3 hidden layers
  64 output dim
  128 hidden dim
  ELU
ResNet
  3 hidden layers
  64 output dim
  256 hidden dim
  ELU
initialization Kaiming Normal Kaiming Normal
AdamW (lr, weight decay) $(5\times 10^{-3},0.0)$ $(5\times 10^{-3},10^{-12})$
StepLR (gamma, step size) $(0.5,50)$ $(0.5,50)$
Table 10: Training details for DE-DeepONet
Dataset
Hyperelasticity Navier–Stokes
branch net ResNet
  3 hidden layers
  128 output dim
  128 hidden dim
  ELU
ResNet
  3 hidden layers
  256 output dim
  256 hidden dim
  ELU
trunk net ResNet
  3 hidden layers
  256 output dim
  256 hidden dim
  ReLU
ResNet
  3 hidden layers
  512 output dim
  512 hidden dim
  ReLU
initialization Kaiming Uniform Kaiming Uniform
AdamW (lr, weight decay) $(10^{-3},10^{-11})$ $(10^{-3},10^{-11})$
disable lr scheduler True True
number of Fourier features $64$ $64$
Fourier feature scale $\sigma$ $0.5$ $1.0$
$N_{x}^{\text{batch}}(=\alpha N_{x})$ $422\approx(0.1\times 65^{2})$ $422\approx(0.1\times 65^{2})$
Table 11: Number of trainable parameters in each model
Dataset # parameters
DeepONet FNO & DE-FNO DINO DE-DeepONet
Hyperelasticity 4.21 M 17.88 M 0.04 M 0.27 M
Navier–Stokes 4.21 M 17.88 M 0.15 M 1.02 M
B.5 Convergence plot
Based on the training time (seconds/iteration) of each model in Table 4, we obtain the convergence plots when the number of training samples is either limited ($N_{\text{train}}=16,64$) or sufficient ($N_{\text{train}}=256,1024$) in Figures 6, 7, 8 and 9 for the hyperelasticity equation, and in Figures 2, 10 and 11 for the Navier–Stokes equations.
@@ B.5 Convergence plot
Figure 6: Mean relative errors ($\pm$ standard deviation) over 5 trials versus model training time for the hyperelasticity equations when the number of training samples is 16.
@@
Figure 7: Mean relative errors ($\pm$ standard deviation) over 5 trials versus model training time for the hyperelasticity equations when the number of training samples is 64.
@@
Figure 8: Mean relative errors ($\pm$ standard deviation) over 5 trials versus model training time for the hyperelasticity equations when the number of training samples is 256.
@@
Figure 9: Mean relative errors ($\pm$ standard deviation) over 5 trials versus model training time for the hyperelasticity equations when the number of training samples is 1024.
@@
Figure 10: Mean relative errors ($\pm$ standard deviation) over 5 trials versus model training time for the Navier–Stokes equations when the number of training samples is 64.
@@
Figure 11: Mean relative errors ($\pm$ standard deviation) over 5 trials versus model training time for the Navier–Stokes equations when the number of training samples is 1024.
B.6 Output reconstruction error